On January 1, 2021, Julee Enterprises borrows $34,000 to purchase a new Toyota Highlander by agreeing to a 6%, 4-year note with the bank. Payments of $798.49 are due at the end of each month with the first installment due on January 31, 2021.

Record the issuance of the note payable and the first two monthly payments. (If no entry is required for a particular transaction/event, select "No Journal Entry Required" in the first account field. Do not round intermediate calculations. Round your answers to 2 decimal places.)

Record the issuance of the note payable on Jan 01, 2021.

Record the first payment on Jan 31, 2021.

Record the second payment on Feb 28, 2021.
Answer #1
Solution:
Date Account titles Debit Credit
Jan'1, 2021 Cash $     34,000
   6% Notes Payable $     34,000
Jan'31, 2021 6% Notes Payable (798.49 - 170) $     628.49
Interest Expenses (34,000*6%*1/12) $     170.00
    Cash $     798.49
Feb'28, 2021 6% Notes Payable $     631.63
Interest Expenses (refer note 1) $     166.86
    Cash $     798.49
Note 1---- Interest Expenses for 2nd monthly payment
Loan Fund Outstanding at the beginning of 2nd month = 34,000 - 628.49 = $33,371.51
Interest Expenses = 33,371.51*6%*1/12 = $166.86